OpenStudy (anonymous):

Solve the given differential equation:

OpenStudy (anonymous):

\[yy''+(y')^2=0\] I think I'm suppose to use some sort of v-sub or using variation of parameters to solve it, but I'm not exactly sure how. If v=y', then v'=y''. But what is y?

OpenStudy (anonymous):

I think I have to change the independent variable? Right now it's in terms of t, but once I make the v-sub then I could make it in terms of y, so y becomes the independent?

OpenStudy (freckles):

\[(y' y)'=y'' y+y' y' =y''y+(y')^2\]

OpenStudy (freckles):

\[y y''+(y')^2=0 \\ (y' y)'=0\]

OpenStudy (anonymous):

Hmm, but what about in the case of y''+y=0 or y''+y(y')^3=0?

OpenStudy (anonymous):

I don't necessarily need those answers, but I think there's a different approach besides inspection to see that the combination is the product rule =/ Because that's not always the case

OpenStudy (freckles):

I know it isn't always the case but it worked here :p

ganeshie8 (ganeshie8):

y'' + y = 0 is a regular linear equation which you can solve by using characteristic eqn thingy

OpenStudy (anonymous):

Okay, but it wants me to solve using the substitution technique XD I should've made that more clear. I just wasn't sure how to go about doing that exactly

OpenStudy (freckles):

use the sub u=y' y :p

OpenStudy (freckles):

\[u=y' y \\ u'=y''y +(y')^2 \\ \text{ so we have } u'=0\]

OpenStudy (anonymous):

That's still by inspection though :P I managed to find a similar solution, and their approach was that \[v=y', v'=y'', \frac{ dv }{ dt }=v \frac{ dy }{ dt }\] But I don't exactly understand this either..

OpenStudy (freckles):

also please don't get mad at me I used to have blonde hair

ganeshie8 (ganeshie8):

In each case, you may let \(\dfrac{dy}{dx}=v\) and get \[y''=v' = \dfrac{dv}{dx} = \dfrac{dv}{dy}\dfrac{dy}{dx} = \dfrac{dv}{dy}v\]

ganeshie8 (ganeshie8):

notice that we're assuming \(v\) as a function of \(y\), not \(x\)

ganeshie8 (ganeshie8):

we're actually gettign rid of the independent variable \(x\), that allows us to work with the differential equaiton with indepent variable \(y\) and dependent variable \(v\)

OpenStudy (anonymous):

@freckles lol I'm not mad :P Okay, that makes a little bit more sense!

ganeshie8 (ganeshie8):

for example, \(y''+y=0\) changes to the system of first order eqns \(\dfrac{dy}{dx}= v(y)\) \(\dfrac{dv}{dy}v(y)+y=0\)

ganeshie8 (ganeshie8):

as you can see, the equation is trivial to solve now because it is separable : \(\dfrac{dv}{dy}v(y)+y=0\) \(\int v(y) dv = \int -y dy\)

ganeshie8 (ganeshie8):

once you have solved \(v(y)\), you can solve \(y(x)\) using the other eqn : \(\dfrac{dy}{dx}=v(y)\)

OpenStudy (anonymous):

Ahh, that makes sense now. Thanks!!

ganeshie8 (ganeshie8):

np